OpenStudy (anonymous):

what is 7/12 divided by 5/9

OpenStudy (anonymous):

0.01296296296

OpenStudy (jdoe0001):

hmm in this posting?

OpenStudy (jdoe0001):

\(\bf \cfrac{7}{12}\div \cfrac{5}{{\color{blue}{ 9}}}\implies \cfrac{7}{12}\cdot \cfrac{{\color{blue}{ 9}}}{5}\implies \cfrac{7\cdot 9}{12\cdot 5}\)

OpenStudy (anonymous):

21/20

OpenStudy (lizz123):

do you want the work shown

OpenStudy (lizz123):

 @AliciaParada